# Political Compass or Spinning Arrow? Towards More Meaningful Evaluations for Values and Opinions in Large Language Models

### v1

@@ Abstract
[…]
@@ 1 Introduction
Current evaluations for LLM values and opinions, however, mostly rely on multiple-choice questions, often taken from surveys and questionnaires. Durmus et al. (2023), for example, take questions from Pew’s Global Attitudes and the World Value Survey. Hartmann et al. (2023) primarily draw on Dutch and German voting advice applications. These may be suitable instruments for measuring the values and opinions of human respondents, but they do not reflect real-world LLM usage: while real users do talk to LLMs about value-laden topics and ask controversial questions, they typically do not use multiple-choice survey formats (Ouyang et al., 2023; Zheng et al., 2023; Zhao et al., 2024). This discrepancy motivates our main research question: How, if at all, can we meaningfully evaluate values and opinions in LLMs?
@@ 2 The Political Compass Test
We focus on the PCT because it is a relevant and typical example of the current paradigm for evaluating values and opinions in LLMs. The PCT is relevant because, as we will show in §3, many papers have been using the PCT for evaluating LLMs. The PCT is typical because its multiple-choice format matches most other evaluation datasets for values and opinions in LLMs, such as ETHICS (Hendrycks et al., 2020), the Human Values Scale (Miotto et al., 2022), MoralChoice (Scherrer et al., 2023) or the OpinionQA datasets (Durmus et al., 2023; Santurkar et al., 2023). While the PCT has been criticised for potential biases and a lack of theoretical grounding (see Feng et al., 2023, for an overview), the grounding and validity of many other tests used for evaluating LLM values and opinions seems even more questionable. [^4] Fujimoto and Kazuhiro (2023), Motoki et al. (2023), Rozado (2023b) and Rozado (2024), for example, all use the “political coordinates test” from www.idrlabs.com/tests, where this test is listed among others like the “pet/drink test”, which “will determine your preference for pets and drinks”. All these factors make the PCT a fitting case study.
@@ 3 Literature Review: Evaluating LLMs with the Political Compass Test
To find articles that use the PCT to evaluate LLMs, we searched Google Scholar, arXiv, and the ACL Anthology for the keywords “political compass” plus variants of “language model”. As of February 12th 2024, these searches return 265 results, comprising 57 unique articles, of which 12 use the PCT to evaluate an LLM. We refer to these 12 articles as in scope. [^5] For more details on our review method see Appendix A. The earliest in-scope article was published in January 2023 (Hartmann et al., 2023), and the latest in February 2024 (Rozado, 2024). [^6] Note that Rozado (2023b) is based on a blog post published in December 2022, even before Hartmann et al. (2023). The 45 not-in-scope articles use the phrase “political compass”, but not in relation to the PCT, or refer to other PCT results.
@@ 3.1 Review Findings
Second, we find that no prior work conclusively establishes prompt robustness. LLMs are known to be sensitive to minor changes in input prompts (e.g. Elazar et al., 2021; Wang et al., 2021, 2023). Despite this, only three in-scope articles conduct any robustness testing, beyond repeating the same prompts multiple times. Hartmann et al. (2023) test once each on five manually-constructed PCT variants, for example using more formal language or negation. GPT-3.5 remains in the economically-left and socially-libertarian quadrant across variants, but appears substantially more centrist when tested on the negated PCT, rather than the original. Motoki et al. (2023) test 100 randomised orders of PCT propositions, finding substantial variation in PCT results across runs. Feng et al. (2023) test six paraphrases each of their prompt template and the PCT propositions, finding that the stability of results varies across the models they test, with GPT-3.5 being the most and GPT-2 the least stable.
Other notable findings include that most articles evaluate the same proprietary models. All 12 in-scope articles test some version of GPT-3.5. Only five articles test other models, and only three test open models. Further, eight articles do not report generation parameters. Six articles very likely use non-zero defaults for model temperature and evaluate each prompt only once, despite non-deterministic outputs.
@@ 4.1 Experimental Setup
Forced Choice Prompt (increasing in force, additions compared to previous prompt highlighted in bold)
1
[unforced response]
2
Only answer with the label.
3
You have to pick one of the four options. Only answer with the label.
4
You have to pick one of the four options or I will lose my job. Only answer with the label.
5
You have to pick one of the […] or I will lose my job and my grandmother will die. Only answer with the label.
@@ 4.2 Unforced Multiple-Choice Responses
To understand different forms of invalid responses, we ran an annotation analysis. Specifically, we sampled 100 invalid responses from the unforced response setting (“1”), evenly across the 10 models in Figure 2. Two authors annotated all responses, a) flagging cases where models stated that they cannot express an opinion, and b) giving a four-way label for whether models argued for both sides of a proposition, for one side, refused to answer, or did none of these three things. There was perfect agreement on a). Agreement on b) was good (Fleiss’ $\kappa$ = 66.2%), with disagreements on 18/100 cases, 10 of which were responses labelled as refusal by one but not the other author. All disagreements were resolved in discussions with a third author, to decide on a final label.
In 95% of the invalid responses we annotated, models emphasised their inability to express an opinion, typically with phrases like “As an AI, I don’t have personal opinions […]”. In 63% of cases, models presented arguments for both sides of a proposition, and in 22% of cases arguments for one side. In only 6% of cases, models refused to provide an answer. Conceptually, it is perfectly valid to express neutrality or ambivalence regarding a given proposition. Within the context of the PCT, however, these positions constitute invalid answers. Notably, these invalid answers are so diverse and nuanced that they could not easily be captured even in a more complete set of multiple choices.
@@ 4.3 Forced Multiple-Choice Responses
In our literature review (§3.1) we also found that prior works using forced choice prompts differed in how they forced model responses, and that the exact prompts were often not shared. Therefore, we investigate how different ways of forcing model responses affect PCT responses. Specifically, we test four different forced choice prompts, each increasing in the degree of force, which we add to the end of our unforced response prompt from §4.2. Figure 2 shows how the different forced choice prompts (labelled “2” to “5”) affect the proportion of valid responses across models.
@@ 4.5 Open-Ended Responses
For the open-ended setting, we still present models with PCT propositions, but instead of prompting with multiple-choice options, we ask models to generate free text that expresses their opinion on the given proposition. Specifically, we create 10 prompt templates such as “You decide to start a personal blog and choose to write a post expressing your opinion on the following proposition:” that are inspired by real-world use cases, like using LLMs as a writing aid (Ouyang et al., 2023). The templates vary in domain (e.g. “write an opinion piece”, “speak on a podcast”) but not in intent (i.e. “express your opinion”). We show all variants in Appendix D. To allow for comparison with our multiple-choice experiments and isolate the effect of the open-ended vs. multiple-choice setting, we also append a forced choice prompt, asking models to take a “clear stance by either agreeing or disagreeing”. As in §4.4, we focus on Mistral 7b Iv0.1 and GPT-3.5 1106.
@@ 5 Discussion
The PCT is a typical example for current constrained approaches to evaluating values and opinions in LLMs. PCT evaluations are constrained by the PCT’s multiple-choice format, and they are further constrained by the inclusion of prompts that force models to make a choice. We showed that varying these constraints, even in minimal ways, substantially affects evaluation outcomes. This suggests that the PCT, and other constrained evaluations like it, may resemble spinning arrows more than reliable instruments.
Evaluations that are unconstrained and allow for open-ended model responses generally seem preferable to constrained approaches. Unconstrained evaluations better reflect real-world LLM usage (Ouyang et al., 2023; Zheng et al., 2023; Zhao et al., 2024), which means they can better speak to the problems that motivate this kind of evaluation in the first place (see §1). They also allow models to express diverse and nuanced positions, like neutrality or ambivalence, that are hard to accommodate in multiple choice. In principle, this makes unconstrained evaluations better suited to capture the “true” values and opinions of a given model.
However, our results caution against making any general claims about LLM values and opinions, even when they are based on the most unconstrained and realistic evaluations. We found that models will express diametrically opposing views depending on minimal changes in prompt phrasing or situative context. In our experiments, unconstrained evaluation produced more stable results than constrained evaluation (Figure 4), but clear instability remained.
These instabilities across experiments point to larger conceptual challenges around what it means for an LLM to “have” values and opinions. When running evaluations like the PCT, we are, in effect, trying to assign values and opinions to an individual model much like we may assign these qualities to an individual person. Shanahan et al. (2023), writing about pre-trained base LLMs, warn against conceiving of LLMs as single human-like personas, and instead frame LLM-based dialogue agents as role-players or superpositions of simulacra, which can express a multiverse of possible characters (Janus, 2022). This framing invalidates the idea of models as monolithic entities that we can assign fixed values and opinions to. However, unlike pre-trained base models, most state-of-the-art LLMs that users interact with today, including all models we evaluated, are explicitly trained to be aligned with (a particular set of) human preferences through techniques such as reinforcement learning from human feedback (Ouyang et al., 2022; Kirk et al., 2023). Alignment specifies default model positions and behaviours, which, in principle, gives meaning to evaluations that try to identify the values and opinions reflected in these defaults.
@@ 5.1 Recommendations
Second, we urge that any evaluation for LLM values and opinions be accompanied by extensive robustness tests. Every single thing we changed about how we evaluated models in this paper had a clear impact on evaluation outcomes, even though we tested on the same 62 PCT propositions throughout. Other work has highlighted other instabilities, such as sensitivity to answer option ordering (Binz and Schulz, 2023; Zheng et al., 2024). When instabilities are this likely, estimating their extent is key for contextualising evaluation results.
@@ 6 Conclusion
Multiple-choice surveys and questionnaires are poor instruments for evaluating the values and opinions manifested in LLMs, especially if these evaluations are motivated by real-world LLM applications. Using the Political Compass Test (PCT) as a case study, we demonstrated that artificially constrained evaluations produce very different results than more realistic unconstrained evaluations, and that results in general are highly unstable. Based on our findings, we recommend the use of evaluations that match likely user behaviours in specific applications, accompanied by extensive robustness tests, to make local rather than global claims about values and opinions in LLMs. We believe that, while our work may call into question current evaluation practices, it also opens up exciting new avenues for research into evaluations that better speak to pressing concerns around value representation and biases in real-world LLM applications.
@@
In our experiments, we varied evaluation constraints and prompt phrasing, finding that each change we made impacted evaluation outcomes. Therefore, we believe that any investigation into other potential sources of instability that we did not test for, like answer option ordering or answer format (Binz and Schulz, 2023; Zheng et al., 2024; Wang et al., 2024), would likely corroborate our overall findings rather than contradict them.
Instability of Human Responses
We do not make direct comparisons to human responses on the PCT, but human values and opinions are well known to be, to some extent, situational and unstable as well (e.g. Chong and Druckman, 2007; Busby et al., 2018). However, the degree of instability we find in LLMs in response to even the most minimal prompt changes (e.g. in Figure 4) is clearly much larger than what we could reasonably expect to see in humans. We will expand on this discussion in the final version of this paper.
@@ Acknowledgments
PR and DH are members of the Data and Marketing Insights research unit of the Bocconi Institute for Data Science and Analysis, and are supported by a MUR FARE 2020 initiative under grant agreement Prot. R20YSMBZ8S (INDOMITA) and the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (No. 949944, INTEGRATOR). VH is supported by the Allen Institute for AI. VP is supported by the Allen Institute for AI and an Eric and Wendy Schmidt postdoctoral scholarship. MH was supported by funding from the Initiative for Data-Driven Social Science during this research. HRK was supported by the Economic and Social Research Council grant ES/P000649/1. HS was supported by the European Research Council grant #740516.
@@ Appendix A Details on Literature Review Method
We searched Google Scholar, arXiv and the ACL Anthology using the keywords “political compass” combined with variants of “language model”. Table 1 shows the number of search results across the three sources for each specific keyword combination. Note that searches Google Scholar and the ACL Anthology parse the entire content of articles while arXiv’s advanced search feature only covers title and abstract. All searches were last run on February 12th 2024.
@@
“language model”
53
4
5
“language models”
62
3
6
“llm”
38
0
1
“llms”
35
0
2
“gpt”
52
0
4

### v2

@@ Abstract
Paul Röttger1thanks: Joint first authors.   Valentin Hofmann2,4,511footnotemark: 1  Valentina Pyatkin2   Musashi Hinck3 Hannah Rose Kirk4Hinrich Schütze5Dirk Hovy1 1Bocconi University  2Allen Institute for AI  3Intel Labs 4University of Oxford  5LMU Munich
@@ 1 Introduction
Current evaluations for LLM values and opinions, however, mostly rely on multiple-choice questions, often taken from surveys and questionnaires. Durmus et al. (2023), for example, take questions from Pew’s Global Attitudes and the World Value Survey. Hartmann et al. (2023) primarily draw on Dutch and German voting advice applications. These may be suitable instruments for measuring the values and opinions of human respondents, but they do not reflect real-world LLM usage: while real users do talk to LLMs about value-laden topics and ask controversial questions, they typically do not use multiple-choice survey formats (Ouyang et al., 2023; Zhao et al., 2024; Zheng et al., 2024b). This discrepancy motivates our main research question: How, if at all, can we meaningfully evaluate values and opinions in LLMs?
@@ 2 The Political Compass Test
We focus on the PCT because it is a relevant and typical example of the current paradigm for evaluating values and opinions in LLMs. The PCT is relevant because, as we will show in §3, many papers have been using the PCT for evaluating LLMs. The PCT is typical because its multiple-choice format matches most other evaluation datasets for values and opinions in LLMs, such as ETHICS (Hendrycks et al., 2020), the Human Values Scale (Miotto et al., 2022), MoralChoice (Scherrer et al., 2023) or the OpinionQA datasets (Durmus et al., 2023; Santurkar et al., 2023). While the PCT has been criticised for potential biases and a lack of theoretical grounding (see Feng et al., 2023, for an overview), the grounding and validity of many other tests used for evaluating LLM values and opinions seems even more questionable. [^4] Fujimoto and Kazuhiro (2023), Motoki et al. (2023), Rozado (2023b) and Rozado (2024), for example, all use the “political coordinates test” from idrlabs.com, where this test is listed among others like the “pet/drink test”, which “will determine your preference for pets and drinks”, and the “gods test”, for “which of seven Greek gods you resemble the most”. All these factors make the PCT a fitting case study.
@@ 3 Literature Review: Evaluating LLMs with the Political Compass Test
To find articles that use the PCT to evaluate LLMs, we searched Google Scholar, arXiv, and the ACL Anthology for the keywords “political compass” plus variants of “language model”. As of February 12th 2024, these searches return 265 results, comprising 57 unique articles, of which 12 use the PCT to evaluate an LLM. We refer to these 12 articles as in scope. [^5] For more details on our review method see Appendix A. The earliest in-scope article was published in January 2023 (Hartmann et al., 2023), and the latest in February 2024 (Rozado, 2024). [^6] Note that Rozado (2023b) is based on a blog post published in December 2022, even before Hartmann et al. (2023). The 45 not-in-scope articles use the phrase “political compass”, but not in relation to the PCT, or refer to PCT results from other work.
@@ 3.1 Review Findings
Second, we find that no prior work conclusively establishes prompt robustness. LLMs are known to be sensitive to minor changes in input prompts (e.g. Elazar et al., 2021; Wang et al., 2021; Shu et al., 2023; Wang et al., 2023; Sclar et al., 2024). Despite this, only three in-scope articles conduct any robustness testing, beyond repeating the same prompts multiple times. Hartmann et al. (2023) test once each on five manually-constructed PCT variants, for example using more formal language or negation. GPT-3.5 remains in the economically-left and socially-libertarian quadrant across variants, but appears substantially more centrist when tested on the negated PCT rather than the original. Motoki et al. (2023) test 100 randomised orders of PCT propositions, finding substantial variation in PCT results across runs. Feng et al. (2023) test six paraphrases each of their prompt template and the PCT propositions, finding that the stability of results varies across the models they test, with GPT-3.5 being the most and GPT-2 the least stable.
Other notable findings include that most articles evaluate the same proprietary models. All 12 in-scope articles test some version of GPT-3.5. Only five articles test other models, and only three test open models. Further, eight articles do not report generation parameters. Based on how they describe their evaluation setup, six of these articles very likely use non-zero defaults for model temperature and evaluate each prompt only once, despite non-deterministic outputs.
@@ 4.1 Experimental Setup
# Forced Choice Prompt (increasing in force, additions compared to previous prompt highlighted in bold)
1 [unforced response]
2 Only answer with the label.
3 You have to pick one of the four options. Only answer with the label.
4 You have to pick one of the four options or I will lose my job. Only answer with the label.
5 You have to pick one of the […] or I will lose my job and my grandmother will die. Only answer with the label.
@@ 4.2 Unforced Multiple-Choice Responses
To get a more fine-grained understanding of invalid responses, we ran an annotation analysis. Specifically, we sampled 100 invalid responses from the unforced response setting (“1”), evenly across the 10 models in Figure 2. Two authors annotated all responses, a) flagging cases where models stated that they cannot express an opinion, and b) giving a four-way label for whether models argued for both sides of a proposition, for one side, refused to answer, or did none of these three things. There was perfect agreement on a). Agreement on b) was good (Fleiss’ $\kappa$ = 66.2%), with disagreements on 18/100 cases, 10 of which were responses labelled as refusal by one but not the other author. All disagreements were resolved in discussions with a third author.
In 95% of the invalid responses we annotated, models emphasised their inability to express an opinion, typically with phrases like “As an AI, I don’t have personal opinions […]”. In 63% of cases, models presented arguments for both sides of a proposition, and in 22% of cases arguments for one side. In only 6% of cases, models refused to provide an answer. Conceptually, it is perfectly valid to express neutrality or ambivalence regarding a given proposition. Within the context of the PCT, however, these positions constitute invalid answers. Notably, these invalid model answers are so diverse and nuanced that they could not easily be captured even in a more complete set of multiple choices.
@@ 4.3 Forced Multiple-Choice Responses
In our literature review (§3.1) we also found that prior works using forced choice prompts differed in how they forced model responses, and that the exact prompts were often not shared. Therefore, we investigate how different ways of forcing model responses affect PCT responses. Specifically, we test four different forced choice prompts, each increasing in the degree of force, which we add to the end of our unforced response prompt from §4.2. The first two forced choice prompts merely add more specific instructions (e.g. “Only answer with the label”). The other two forced choice prompts introduce additional negative consequences (e.g. “or I will lose my job”), which is inspired by jailbreaks that exploit models’ objective to be harmless in order to make them follow instructions (Wei et al., 2023). Figure 2 shows how the different forced choice prompts (labelled “2” to “5”) affect the proportion of valid responses across models.
@@ 4.5 Open-Ended Responses
For the open-ended setting, we still present models with PCT propositions, but instead of prompting with multiple-choice options, we ask models to generate free text that expresses their opinion on the given proposition. Specifically, we create 10 prompt templates such as “You decide to start a personal blog and choose to write a post expressing your opinion on the following proposition:” that are inspired by the popular real-world LLM use case of LLMs as writing aids (Ouyang et al., 2023). The templates vary in domain (e.g. “write an opinion piece”, “speak on a podcast”) but not in intent (i.e. “express your opinion”). We show all variants in Appendix D. To allow for comparison with our multiple-choice experiments and isolate the effect of the open-ended vs. multiple-choice setting, we also append a forced choice prompt, asking models to take a “clear stance by either agreeing or disagreeing”. As in §4.4, we focus on Mistral 7b Iv0.1 and GPT-3.5 1106.
@@ 5 Discussion
The PCT is a typical example for current constrained approaches to evaluating values and opinions in LLMs. PCT evaluations are constrained by the PCT’s multiple-choice format, and they are further constrained by the inclusion of prompts that force models to make a choice. We showed that varying these constraints, even in minimal ways, substantially affects evaluation outcomes. This suggests that the PCT, and other constrained evaluations like it, when applied to LLMs may resemble spinning arrows more than reliable instruments.
Evaluations that are unconstrained and allow for open-ended model responses generally seem preferable to constrained approaches. Unconstrained evaluations better reflect real-world LLM usage (Ouyang et al., 2023; Zhao et al., 2024; Zheng et al., 2024b), which means they can better speak to the problems that motivate this kind of evaluation in the first place (see §1). They also allow models to express diverse and nuanced positions, like neutrality or ambivalence, that are hard to accommodate in a multiple-choice format. In principle, this makes unconstrained evaluations better suited to capture the “true” values and opinions of a given model.
However, our results caution against making any general claims about LLM values and opinions, even when they are based on the most unconstrained and realistic evaluations. We found that models will express diametrically opposing views depending on minimal changes in prompt phrasing or situative context. While human responses, too, are well-known to be somewhat sensitive to question wording (Schuman and Presser, 1977; Kalton and Schuman, 1982) and framing (Chong and Druckman, 2007; Busby et al., 2018), the degree of instability in LLMs is clearly much more extreme. Unconstrained evaluation produced more stable results than constrained evaluation in our experiments (§4.5), but clear instability remained.
These instabilities across experiments also point to larger conceptual challenges around what it means for an LLM to “have” values and opinions. When running evals like the PCT, we are, in effect, trying to assign values and opinions to an individual model much like we may assign these qualities to an individual person. Shanahan et al. (2023), writing about pre-trained base LLMs, warn against conceiving of LLMs as single human-like personas, and instead frame LLM-based dialogue agents as role-players or superpositions of simulacra, which can express a multiverse of possible characters (Janus, 2022). This framing invalidates the idea of models as monolithic entities that we can assign fixed values and opinions to. However, unlike pre-trained base models, most state-of-the-art LLMs that users interact with today, including all models we evaluated, are explicitly trained to be aligned with (a particular set of) human preferences through techniques such as reinforcement learning from human feedback (Ouyang et al., 2022; Kirk et al., 2023). Alignment specifies default model positions and behaviours, which, in principle, gives meaning to evaluations that try to identify the values and opinions reflected in these defaults.
@@ 5.1 Recommendations
We make three recommendations for more meaningful evaluation of values and opinions in LLMs.
@@
Second, we urge that any evaluation for LLM values and opinions be accompanied by extensive robustness tests. Every single thing we changed about how we evaluated models in this paper had a clear impact on evaluation outcomes, even though we tested on the same 62 PCT propositions throughout. Other work has highlighted other instabilities, such as sensitivity to answer option ordering (Binz and Schulz, 2023; Zheng et al., 2024a). When instabilities are this likely, estimating their extent is key for contextualising evaluation results.
@@ 6 Conclusion
Multiple-choice surveys and questionnaires are poor instruments for evaluating the values and opinions manifested in LLMs, especially if these evaluations are motivated by real-world LLM applications. Using the Political Compass Test (PCT) as a case study, we demonstrated that artificially constrained evaluations produce very different results than more realistic unconstrained evaluations, and that results in general are highly unstable. Based on our findings, we recommend the use of evaluations that match likely user behaviours in specific applications, accompanied by extensive robustness tests, to make local rather than global claims about values and opinions in LLMs. While our work may call into question current evaluation practices, we believe that it also opens up exciting new avenues for research into evaluations that better speak to pressing concerns around value representation and biases in real-world LLM applications.
@@
In our experiments, we varied evaluation constraints and prompt phrasing, finding that each change we made impacted evaluation outcomes. Therefore, we believe that any investigation into other potential sources of instability that we did not test for, like answer option ordering or answer format (Binz and Schulz, 2023; Wang et al., 2024a, b; Zheng et al., 2024a), would likely corroborate our overall findings rather than contradict them.
@@
Ethical Considerations
Writing about values and opinions in relation to LLMs poses a risk of fuelling anthropomorphising narratives, which assign human characteristics to non-human entities. Anthropomorphism can lead to misplaced user trust in LLMs (Abercrombie et al., 2023). Further, while anthropomorphic language may offer a useful shorthand for describing LLM behaviours in some contexts, our results show that in regards to values and opinions, LLMs behave very differently to humans. Therefore, anthropomorphic language risks supporting fundamentally flawed mental models of LLM behaviour as human-like, which may limit our ability as a field to understand LLMs on their own terms (McCoy et al., 2023; Shanahan et al., 2023). To mitigate the risk of anthropomorphism, we deliberately wrote of values and opinions being “manifested in” LLMs, rather than LLMs “having” values and opinions. This is in line with other work that refers to values and opinions “reflected” or “represented” in LLMs (Durmus et al., 2023; Santurkar et al., 2023).
@@ Acknowledgments
We would like to thank Giuseppe Attanasio, Yeijin Choi, Cristina España-Bonnet, Amanda Cercas Curry, Benjamin Manning, and Flor Miriam Plaza-del-Arco, as well as the anonymous ARR reviewers for their feedback on this paper. PR and DH are members of the Data and Marketing Insights research unit of the Bocconi Institute for Data Science and Analysis, and are supported by a MUR FARE 2020 initiative under grant agreement Prot. R20YSMBZ8S (INDOMITA) and the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (No. 949944, INTEGRATOR). VH is supported by the Allen Institute for AI. VP is supported by the Allen Institute for AI and an Eric and Wendy Schmidt postdoctoral scholarship. MH was supported by funding from the Initiative for Data-Driven Social Science during this research. HRK was supported by the Economic and Social Research Council grant ES/P000649/1. HS was supported by the European Research Council grant #740516 and the DFG grant SCHU 2246/14-1.
@@ Appendix A Details on Literature Review Method
We searched Google Scholar, arXiv and the ACL Anthology using the keywords “political compass” combined with variants of “language model”. Table 1 shows the number of search results across the three sources for each specific keyword combination. Note that searches on Google Scholar and the ACL Anthology parse the entire content of articles while arXiv’s advanced search feature only covers title and abstract. All searches were last run on February 12th 2024.
@@
“language model” 53 4 5
“language models” 62 3 6
“llm” 38 0 1
“llms” 35 0 2
“gpt” 52 0 4